# Enhancing Hemodynamic Parameter Estimations: Nonlinear Blood Behavior in 4D Flow MRI

### v2

@@ 3.4 Experiment 1: evaluating the pertinence of the power-law
In this experiment, we aimed to assess whether Newtonian and non-Newtonian viscosity models yield equivalent results in terms of WSS, $\dot{E}_{L}$, and OSI across different Hct levels. The goal was to determine if a constant viscosity model suffices to account for the effects of nonlinear viscosity on hemodynamic parameters. If the results between the models are similar, a constant viscosity would be deemed adequate. Otherwise, the adoption of a more complex model would be necessary.
To evaluate this, CFD simulations of a patient-specific aorta model (see Figure 1) were performed for the Hct values listed in Table 1, using both the fitted-Newtonian viscosities ($\mu_{\text{NF}}$) and power-law viscosities ($\mu_{\text{PL}}$) estimated as described in Section 2.4. Details of the model are provided in Section 3.1. The resulting velocity fields were then used to generate synthetic 4D Flow images as described in Section 2.3, which were processed to obtain WSS, $\dot{E}_{L}$, and OSI using both $\mu_{\text{NF}}$ and $\mu_{\text{PL}}$ accordingly. The procedure used to derive these parameters is detailed in Section 3.3.
@@ 3.5 Experiment 2: comparison of current approaches with power-Law model parameter estimation
Given that blood exhibits shear-thinning behavior, this experiment aims to demonstrate the differences in the estimation of hemodynamic parameters when using a Newtonian viscosity ($\mu_{\text{N}}$) compared to a model that accounts for shear-thinning behavior ($\mu_{\text{PL}}$). To illustrate this, the same 4D Flow images generated in Experiment 1 from power-law viscosity CFD simulations were used to obtain WSS, $\dot{E}_{L}$, and OSI at different Hct levels. These parameters were estimated using both Newtonian viscosities within the typical range found in the literature and power-law viscosities as described in Section 2.4. It is important to note that the hemodynamic parameters estimated using the power-law model are exactly the same that those from Experiment 1.
@@ 4 Results
Figure 6: Velocity maps at peak systole and diastole, extracted from CFD simulations using power-law and Newtonian-fitted viscosities for different Hct values. Almost identical flow patterns are observed at peak systole, while some differences are noticeable at diastole, particularly in the descending aorta.
@@ 4.1 Experiment 1
Figure 7: Mean values and standard deviations of WSS, OSI, and $\dot{E}_{L}$ across four aortic segments were estimated from simulated images obtained from CFD simulations using power-law ($\mu_{\mathrm{PL}}$) and fitted-Newtonian ($\mu_{\mathrm{NF}}$) viscosities. The estimations were performed using viscosity models and values consistent with those used in the CFD simulations from which synthetic 4D Flow images were obtained. Clear differences were observed between the models for varying Hct values and cardiac phases, indicating that the two models cannot be used interchangeably.
@@
Figure 6 shows the simulated velocities obtained using power-law ($\mu_{\mathrm{PL}}$) and fitted-Newtonian ($\mu_{\mathrm{NF}}$) viscosities. Both models produced nearly identical velocity profiles at peak systole across all Hct levels, while minor differences were observed in the descending aorta during diastole. Despite these differences, the overall flow dynamics between the two models also appear similar during diastole.
@@
Despite the similarities in flow dynamics between the two models, Figure 7 shows that the hemodynamic parameters estimated from synthetic 4D Flow images exhibit noticeable differences at both systole and diastole. At systole, the average differences in WSS and $\dot{E}_{L}$ across all Hct levels, relative to the power-law model, were $119.3\pm 21.7$% and $61.8\pm 14.4$%, respectively. In absolute terms, these differences range from $0.458$ to $2.883$ Pa in WSS and from $0.089$ to $0.911$ $\mu$W for Hct values of 20 and 70, respectively. In all cases, the sign indicates whether the parameter estimated using fitted-Newtonian ($\mu_{\mathrm{NF}}$) viscosities was greater (positive) or smaller (negative) compared to the power-law estimations.
@@
In diastole, the differences observed in systole show inverted signs. The average differences in WSS and $\dot{E}_{L}$ across all Hct levels were $-14.7\pm 5.4$% and $-43.3\pm 3.7$%, respectively. In absolute terms, these differences range from $-0.008$ to $-0.072$ Pa in WSS and from $-0.015$ to $-0.100$ $\mu$W for Hct values of 20 and 70, respectively. Although the absolute differences are smaller, the relative differences and inverted signs suggest that using a Newtonian model for blood, even when adjusted by Hct level, is not suitable for capturing the nonlinear interactions of blood flow. Therefore, a Newtonian model cannot be used as a surrogate, justifying the inclusion of nonlinear rheology.
@@
For OSI, the average difference across all Hct levels was $-33.4\pm 2.8$%. In absolute terms, the differences ranged from $-0.050$ to $-0.064$ for Hct values of 20 and 70, respectively. Since OSI is a metric calculated over the entire cardiac cycle, the negative sign indicates that the WSS vector experienced greater variation when the blood exhibited nonlinear behavior, consistent with our findings stated in previous paragraphs.
@@
Additionally, greater variability between segments was observed in $\dot{E}_{L}$ at peak systole compared to WSS and OSI (see Figure 7). This is more clearly illustrated in Figure 1 in A, which shows the means and standard deviations of hemodynamic parameters within each aortic segment. Higher values were found in the ascending aorta, primarily due to the flow dynamics (see Figure 6).
Finally, Figure 8 shows the WSS calculated directly from the FE simulations using both models at peak systole and diastole. The same trends in differences are observed in these simulations as those obtained from the synthetic 4D Flow MR images.
@@
Figure 8: WSS maps at peak systole and diastole were estimated from CFD simulations using power-law ($\mu_{\mathrm{PL}}$) and fitted-Newtonian ($\mu_{\mathrm{NF}}$) viscosities for different Hct values. The WSS obtained from the fitted-Newtonian simulations is greater at peak systole but smaller at diastole compared to those obtained using the power-law model, indicating that the two simulations cannot be used interchangeably.
@@ 4.2 Experiment 2
Given that blood exhibits shear-thinning behavior, Figure 9 shows the hemodynamic parameters estimated from synthetic 4D Flow images generated from simulations with a power-law blood rheology. The parameters were estimated using both power-law ($\mu_{\mathrm{PL}}$) and standard Newtonian ($\mu_{\mathrm{N}}$) viscosities commonly found in the literature. At systole, the differences in WSS and $\dot{E}_{L}$ across all Hct levels (relative to the power-law model and using $\mu_{\mathrm{N}}=0.0035~{}\text{Pa}\cdot\text{s}$ as reference) ranged from $189.4\%$ to $-41.3\%$ and from $112.9\%$ to $-60.4\%$ for Hct values of 20 and 70, respectively. In absolute terms, these differences ranged from $0.870$ to $-0.908$ Pa for WSS and from $0.238$ to $-0.686$ $\mu$W for Hct values of 20 and 70, respectively. The sign change indicates that WSS and $\dot{E}_{L}$ estimated using Newtonian viscosities can be either greater (positive) or smaller (negative) than the power-law estimations, depending on the Hct level. These differences can be better observed in Figure 10 for WSS maps obtained on the segmented aorta at peak systole.
@@
In diastole, the results were similar, with relative differences in WSS ranging from $65.5\%$ to $-71.8\%$ (using $\mu_{\mathrm{N}}=0.0035~{}\text{Pa}\cdot\text{s}$ as reference) and from $54.9\%$ to $-74.0\%$ in $\dot{E}_{L}$ at peak systole. However, during this cardiac phase, the sign change for WSS occurs at an Hct of 32.5, which is earlier than at peak systole. In absolute terms, these differences ranged from $0.0491$ to $-0.278$ Pa for WSS and from $0.022$ to $-0.153$ $\mu$W for Hct values of 20 and 70, respectively. These differences can be better observed in Figure 11 for WSS maps obtained on the segmented aorta at diastole.
@@
For OSI, higher values were obtained using constant Newtonian viscosities ($\mu_{\mathrm{N}}$) from images generated by power-law simulations, compared to those obtained with fitted-Newtonian viscosities ($\mu_{\mathrm{NF}}$) in Experiment 1 (see Figures 7 and 9). This can be attributed to variations in flow dynamics between the linear and nonlinear rheology models, which are not apparent in the overall flow pattern showed in Figure 6 but significantly affect the velocity gradients near the vessel walls. Using $\mu_{\mathrm{N}}=0.0035~{}\text{Pa}\cdot\text{s}$ as reference, the relative differences with respect to the power-law estimations ranged from $-19.9\%$ to $-27.0\%$ for Hct values of 20 and 70, respectively. In absolute terms, the differences ranged from $-0.033$ to $-0.047$ for the same Hct range. Again, the negative sign indicates that the WSS vector experienced greater temporal variations when the blood exhibited a nonlinear behavior.
Similar to the results in Experiment 2, greater variability between segments was observed in $\dot{E}_{L}$ at peak systole compared to WSS and OSI (see Figure 9). This is more clearly illustrated in Figure 2 in A, which shows the means and standard deviations of hemodynamic parameters within each aortic segment.
Since we have shown differences at peak systole and diastole between estimations made using both viscosity models ($\mu_{\mathrm{N}}$ and $\mu_{\mathrm{PL}}$), it is important to determine whether the parameters estimated using the power-law model are consistent with the simulations. These similarities are highlighted in Figure 8, which shows that the WSS maps estimated from images and CFD simulations exhibit similar spatial variations. However, smaller WSS values were found at peak systole from synthetic 4D Flow MR images compared to their simulation counterparts, with differences around 40% of the CFD simulation values. At diastole, the same WSS values were observed, but the estimations obtained from images were highly affected by noise.
@@ 4.3 Experiment 3
Figure 13: Mean values and standard deviations of WSS, OSI, and $\dot{E}_{L}$ across four aortic segments were estimated from in-vivo 4D Flow images of HCM patients. Hemodynamic parameters were estimated using both power-law ($\mu_{\mathrm{PL}}$) and state-of-the-art Newtonian ($\mu_{\mathrm{N}}$) viscosities. Clear differences were observed between the models across varying Hct values and cardiac phases, indicating that the choice of viscosity model can lead to significant discrepancies in hemodynamic parameter estimations.
@@
Figure 13 shows the hemodynamic parameters estimated from in-vivo 4D Flow MR images. It should be noted that the Hct range in this experiment differs from that in Experiments 1 and 2, ranging from 28.2 to 50.1, compared to 20 to 70 in the previous experiments. Additionally, the Hct values in this case are not equispaced. As evident from the figure, using the power-law viscosity ($\mu_{\mathrm{PL}}$) for estimating hemodynamic parameters resulted in greater values in almost all cases. The only instance where WSS and $\dot{E}_{L}$ were smaller was at an Hct of 28.2 during peak systole.
@@
At peak systole, for an Hct of 28.2 (the lowest), the relative differences with respect to the power-law estimation (using $\mu_{\mathrm{N}}=0.0035~{}\text{Pa}\cdot\text{s}$ as a reference) were $41.4\%$ in WSS and $1.5\%$ in $\dot{E}_{L}$. For an Hct of 50.1 (the highest), the differences were $-21.0\%$ and $-45.4\%$, respectively. However, the largest differences were observed at an Hct of 46.6, with $-45.0\%$ in WSS and $-60.4\%$ in $\dot{E}_{L}$. In absolute terms, these differences in WSS and $\dot{E}_{L}$ were $0.125$ Pa and $0.001$ $\mu$W, $-0.543$ Pa and $-0.101$ $\mu$W, and $-0.312$ Pa and $-0.081$ $\mu$W for Hct levels of 28.2, 50.1, and 46.6, respectively.
@@
In diastole, these differences become larger. For an Hct of 28.2, the relative differences were $-21.4\%$ in WSS and $-34.8\%$ in $\dot{E}_{L}$, while for an Hct of 50.1, they were $-64.6\%$ and $-71.8\%$, respectively. However, as with the peak-systolic results, the largest differences were observed at an Hct of 46.6, with $-69.0\%$ in WSS and $73.0\%$ in $\dot{E}_{L}$. In absolute terms, these differences in WSS and $\dot{E}_{L}$ were $-0.017$ Pa and $-0.005$ $\mu$W, $-0.149$ Pa and $-0.009$ $\mu$W, and $-0.090$ Pa and $-0.016$ $\mu$W for Hct levels of 28.2, 50.1, and 46.6, respectively.
@@
For OSI, the relative differences with respect to the power-law estimations averaged $-21.1\pm 6.5\%$ ($\pm$ one standard deviation) across all Hct values considered in the experiment. In absolute terms, this variation averaged $-0.041\pm 0.019$. These results indicate that using $\mu_{\mathrm{PL}}$ to estimate WSS reveals greater temporal variations, consistent with the findings from Experiment 2.
@@
Although smaller differences between models were found in the in-vivo data, this can be explained by the different blood dynamics and lower flows observed at peak systole compared to the CFD simulations (see Figure 5). The reduced flow likely generates smaller velocity gradients near the vessel walls, leading to lower estimations of WSS and $\dot{E}_{L}$. OSI, on the other hand, is a parameter normalized with respect to time, which helps to maintain the observed differences.
@@
Finally, Figures 14 and 15 show the WSS maps estimated from each patient’s data at peak systole and diastole, respectively. The differences described in the previous paragraphs are clearly illustrated in these figures. Additionally, readers are referred to Figure 3 in A for the means and standard deviations of WSS, $E_{L}$, and OSI within each aortic segment.

### v3

@@ 3.4 Experiment 1: evaluating the pertinence of the power-law
This expermient aims to assess using CFD simulations whether Newtonian and non-Newtonian viscosity models yield equivalent results in terms of WSS, $\dot{E}_{L}$, and OSI across different Hct levels. The goal is to determine if a constant viscosity model suffices to account for the effects of nonlinear viscosity on hemodynamic parameters. If the results between the models are similar, a constant viscosity would be deemed adequate, supporting the current MRI community’s practice of using a constant viscosity for estimating hemodynamic parameters from 4D Flow measurements. Otherwise, the adoption of a more complex model would be necessary.
To evaluate this, CFD simulations of a patient-specific aorta model (see Figure 1) were performed for the Hct values listed in Table 1, using both the Newtonian fitted viscosities ($\mu_{\text{NF}}$) and power-law viscosities ($\mu_{\text{PL}}$) estimated as described in Section 2.4. Details of the model are provided in Section 3.1. The resulting velocity fields were then used to generate synthetic 4D Flow MR images as described in Section 2.3, which were image-processed to obtain WSS, $\dot{E}_{L}$, and OSI using both $\mu_{\text{NF}}$ and $\mu_{\text{PL}}$ with the software and procedure describe in Section 3.3.
@@ 3.5 Experiment 2: comparison of current approaches with power-Law model parameter estimation
Since blood exhibits shear-thinning behavior, this experiment aims to highlight the differences in hemodynamic parameter estimation when using a Newtonian viscosity ($\mu_{\text{N}}$) versus a shear-thinning model ($\mu_{\text{PL}}$). To do so, we used the same 4D Flow images generated in Experiment 1 from power-law viscosity CFD simulations to compute WSS, $\dot{E}_{L}$, and OSI at different Hct levels. These parameters were estimated using both Newtonian (within the typical range found in the literature) and power-law viscosities as described in Section 2.4. Notably, the hemodynamic parameters estimated using $\mu_{\text{PL}}$ are identical to those from Experiment 1.
@@ 4 Results
Figure 7: Velocity maps at peak systole and diastole, extracted from CFD simulations using power-law and Newtonian fitted viscosities for different Hct values. Almost identical flow patterns are observed at peak systole, while great differences are noticeable at diastole.
@@ 4.1 Experiment 1
Figure 8: Mean values and standard deviations of WSS, OSI, and $\dot{E}_{L}$ across four aortic segments were estimated from simulated images obtained from CFD simulations using power-law ($\mu_{\mathrm{PL}}$) and Newtonian fitted ($\mu_{\mathrm{NF}}$) viscosities. The estimations were performed using viscosity models and values consistent with those used in the CFD simulations from which synthetic 4D Flow images were obtained. Clear differences were observed between the models for varying Hct values and cardiac phases, indicating that the two models cannot be used interchangeably.
@@
Figure 7 shows the simulated velocities obtained using power-law ($\mu_{\mathrm{PL}}$) and Newtonian fitted ($\mu_{\mathrm{NF}}$) viscosities. Both models produced nearly identical velocity profiles at peak systole across all Hct levels, while notable differences were observed during diastole. For both models, as the Hct increases (and consequently the viscosity), fluid vorticity decreases at diastole, with this effect being more pronounced in the power-law simulations.
@@
Despite the almost identical velocity maps in systole, Figure 8 shows that hemodynamic parameters estimated from synthetic 4D Flow images exhibit noticeable differences at both systole and diastole. At systole, the average differences in WSS and $\dot{E}_{L}$ across all Hct levels, relative to the power-law model, were $-22.0\pm 1.8$% and $-44.7\pm 3.8$%, respectively. In absolute terms, these differences range from $-0.09$ to $-0.52$ Pa in WSS and from $-0.08$ to $-0.54$ $\mu$W in $\dot{E}_{L}$ for Hct values of 20 and 70, respectively. In all cases, the sign indicates whether the parameter estimated using Newtonian fitted ($\mu_{\mathrm{NF}}$) viscosities was greater (positive) or smaller (negative) compared to the power-law estimations.
@@
In diastole, the relative differences become more pronounced. The average differences in WSS and $\dot{E}_{L}$ across all Hct levels were $-59.9\pm 5.5$% and $-63.3\pm 5.9$%, respectively. In absolute terms, these differences range from $-0.04$ to $-0.23$ Pa in WSS and from $-0.01$ to $-0.09$ $\mu$W in $\dot{E}_{L}$ for Hct values of 20 and 70, respectively. Although the absolute differences are smaller, the relative differences suggest that using a Newtonian model for blood, even when adjusted by Hct level, is not suitable for capturing the nonlinear interactions of blood flow. Consequently, a Newtonian model cannot serve as a suitable surrogate, reinforcing the need to incorporate nonlinear rheology to the estimation of hemodynamic parameters from velocity measurements obtained using 4D Flow MRI.
@@
For OSI, the average difference across all Hct levels was $-24.6\pm 2.6$%. In absolute terms, the differences ranged from $-0.03$ to $-0.05$ for Hct values between 20 and 70, respectively. Since OSI is a metric calculated over the entire cardiac cycle, the negative sign indicates that the WSS vector experienced greater variation when the blood exhibited nonlinear behavior, aligning with our previous findings.
@@
Additionally, greater variability between segments, as indicated by the standard deviation bars, was observed in $\dot{E}_{L}$ at peak systole and OSI compared to WSS (see Figure 8). This is further illustrated in Figure 1 in A, which presents the mean and standard deviation of hemodynamic parameters within each aortic segment. For $\dot{E}_{L}$, higher values were observed in the ascending aorta, with a decreasing trend along the artery, primarily due to the flow dynamics (see Figure 7). A similar pattern is observed for OSI, although the trend is less pronounced in the graph.
Finally, Figure 9 presents the WSS calculated directly from the FE simulations using both models at peak systole and diastole. While differences between models are clearly observable in both MR images and CFD simulations, a notable discrepancy in behavior was observed at peak systole when analyzing the WSS obtained from CFD simulations. In MRI results, WSS estimated using the power-law viscosity model was higher than that obtained with its Newtonian counterpart (see Figure 8). However, in CFD results, the opposite trend was observed, with the Newtonian model yielding higher WSS values than the power-law model. This finding underscores the importance of incorporating a non-Newtonian model, as WSS estimations using the power-law model in MRI exhibited less underestimation compared to its Newtonian counterpart against CFD results.
@@
Figure 9: WSS maps at peak systole and diastole were estimated from CFD simulations using power-law ($\mu_{\mathrm{PL}}$) and Newtonian fitted ($\mu_{\mathrm{NF}}$) viscosities for different Hct values. The color scale for Newtonian fitted results is higher at peak systole. The WSS obtained from the Newtonian fitted simulations is greater at peak systole but smaller at diastole compared to those obtained using the power-law model, indicating that the two simulations cannot be used interchangeably.
@@ 4.2 Experiment 2
Given that blood exhibits shear-thinning behavior (we emphasize that our study builds upon this), Figure 10 shows the hemodynamic parameters estimated from synthetic 4D Flow images generated from simulations with a power-law blood rheology. The parameters were estimated using both power-law ($\mu_{\mathrm{PL}}$) and standard Newtonian ($\mu_{\mathrm{N}}$) viscosities commonly found in the literature. At systole, the differences in WSS and $\dot{E}_{L}$ across all Hct levels (relative to the power-law model and using $\mu_{\mathrm{N}}=0.0035~{}\text{Pa}\cdot\text{s}$ as reference) ranged from $+189.4\%$ to $-41.6\%$ and from $+112.1\%$ to $-60.4\%$ for Hct values of 20 and 70, respectively. In absolute terms, these differences ranged from $+0.87$ to $-0.92$ Pa in WSS and from $+0.23$ to $-0.68$ $\mu$W in $\dot{E}_{L}$ for Hct values from 20 to 70, respectively. The sign change indicates that WSS and $\dot{E}_{L}$ estimated using Newtonian viscosities can be either greater (positive) or smaller (negative) than the power-law estimations, depending on the Hct level. These differences can be better observed in Figure 11 for WSS maps obtained on the segmented aorta at peak systole.
@@
In diastole, the results were similar, with relative differences in WSS ranging from $+51.5\%$ to $-74.4\%$ (using $\mu_{\mathrm{N}}=0.0035~{}\text{Pa}\cdot\text{s}$ as reference) and from $+20.3\%$ to $-79.9\%$ in $\dot{E}_{L}$. In absolute terms, these differences ranged from $+0.03$ to $-0.26$ Pa in WSS and from $+0.01$ to $-0.10$ $\mu$W in $\dot{E}_{L}$ for Hct values from 20 to 70, respectively. In contrast to the systolic results, both WSS and $\dot{E}_{L}$ estimated at diastole using the power-law viscosity model were higher than those obtained with Newtonian viscosities at nearly all Hct levels. These differences are further illustrated in Figure 12, which presents WSS maps of the segmented aorta obtained from 4D Flow MR images at diastole.
@@
For OSI, using $\mu_{\mathrm{N}}=0.0035~{}\text{Pa}\cdot\text{s}$ as the reference, the relative differences with respect to the power-law estimations ranged from $-22.5\%$ to $-30.2\%$ for Hct values between 20 and 70. In absolute terms, the differences ranged from $-0.03$ to $-0.05$ over the same Hct range. As before, the negative sign indicates that the WSS vector exhibited greater temporal variation when blood was modeled with nonlinear behavior.
Similar to the results in Experiment 1, greater variability between segments was observed in $\dot{E}_{L}$ at peak systole and OSI compared to WSS (see Figure 10). This is more clearly illustrated in Figure 2 in A, which shows the means and standard deviations of hemodynamic parameters within each aortic segment.
Since we have demonstrated differences at peak systole and diastole between estimations made using both viscosity models ($\mu_{\mathrm{N}}$ and $\mu_{\mathrm{PL}}$), it is important to assess whether the parameters estimated using the power-law model are consistent with the simulations. These similarities are illustrated in Figure 9, which shows that the WSS maps estimated from MR images and CFD simulations exhibit comparable spatial variations. However, lower WSS values were observed at peak systole in synthetic 4D Flow MR images compared to their simulation counterparts, with differences reaching approximately 60% of the CFD simulation values. At diastole, the WSS values were more comparable, although the estimations obtained from MR images were notably affected by noise.
@@ 4.3 Experiment 3
Figure 14: Mean values and standard deviations of WSS, OSI, and $\dot{E}L$ across four aortic segments were estimated from in-vivo 4D Flow images of HCM patients. Hemodynamic parameters were calculated using both power-law ($\mu{\mathrm{PL}}$) and state-of-the-art Newtonian ($\mu_{\mathrm{N}}$) viscosities. Clear differences were observed between the models across varying Hct values and cardiac phases, highlighting that the choice of viscosity model can result in significant discrepancies in hemodynamic parameter estimations.
@@
Figure 14 presents the hemodynamic parameters estimated from in-vivo 4D Flow MR images. It is important to note that the Hct range in this experiment differs from that used in Experiments 1 and 2, ranging from 28.2 to 50.1, compared to 20 to 70 in the earlier experiments. Additionally, the Hct values in this case are not equidistant. As shown in the figure, estimating hemodynamic parameters using the power-law viscosity model ($\mu_{\mathrm{PL}}$) generally resulted in higher values across nearly all cases. The only exception occurred at an Hct of 28.2 during peak systole, where WSS and $\dot{E}L$ values obtained with $\mu{\mathrm{PL}}$ were lower.
@@
At peak systole, for an Hct of 28.2 (the lowest), the relative differences with respect to the power-law estimation (using $\mu_{\mathrm{N}}=0.0035~{}\text{Pa}\cdot\text{s}$ as a reference) were $+41.4\%$ in WSS and $+1.5\%$ in $\dot{E}_{L}$. For an Hct of 50.1 (the highest), the differences were $-21.0\%$ and $-45.4\%$, respectively. However, the largest differences were observed at an Hct of 46.6, with $-45.0\%$ in WSS and $-60.4\%$ in $\dot{E}_{L}$. In absolute terms, these differences in WSS and $\dot{E}_{L}$ were $+0.125$ Pa and $+0.001$ $\mu$W, $-0.543$ Pa and $-0.101$ $\mu$W, and $-0.312$ Pa and $-0.081$ $\mu$W for Hct levels of 28.2, 50.1, and 46.6, respectively.
@@
In diastole, these differences become larger. For an Hct of 28.2, the relative differences were $-21.4\%$ in WSS and $-34.8\%$ in $\dot{E}_{L}$, while for an Hct of 50.1, they were $-64.6\%$ and $-71.8\%$, respectively. However, as with the peak-systolic results, the largest differences were observed at an Hct of 46.6, with $-69.0\%$ in WSS and $+73.0\%$ in $\dot{E}_{L}$. In absolute terms, these differences in WSS and $\dot{E}_{L}$ were $-0.017$ Pa and $-0.005$ $\mu$W, $-0.149$ Pa and $-0.009$ $\mu$W, and $-0.090$ Pa and $-0.016$ $\mu$W for Hct levels of 28.2, 50.1, and 46.6, respectively.
@@
For OSI, the relative differences with respect to the power-law estimations averaged $-21.1\pm 6.5\%$ across all Hct values considered in the experiment. In absolute terms, this variation averaged $-0.041\pm 0.019$. These results indicate that using $\mu_{\mathrm{PL}}$ to estimate WSS reveals greater temporal variations, consistent with the findings from Experiment 2.
@@
Although smaller differences between viscosity models were observed in the in-vivo MR data (particularly at peak systole), these discrepancies can be attributed to the differing blood flow dynamics and lower flow rates found in patients compared to the CFD simulations (see Figure 6). The reduced flow likely results in smaller velocity gradients near the vessel walls, leading to lower estimations of WSS and $\dot{E}_{L}$. OSI, however, is a parameter normalized over time, which helps preserve the observed differences between models. Additionally, these smaller differences are accentuated by the narrower Hct range considered in this experiment.
@@
Finally, Figures 15 and 16 show the WSS maps estimated from each patient’s data at peak systole and diastole, respectively. The differences described in the previous paragraphs are clearly illustrated in these figures. Additionally, readers are referred to Figure 3 in A for the means and standard deviations of WSS, $E_{L}$, and OSI within each aortic segment.